Find the critical numbers of the function. (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.

1a. f(x) = 48x − 4x2

1b. g(x) = x4 − 2x2 + 2

1c.y =

x
x2 + 49

1d. f(x) = 9 − 9x, Find the open intervals on which the function is increasing or decreasing. Use a graphing utility to verify your results. (Enter your answers using interval notation. If an answer does not exist, enter DNE.)

1e. f(x) = −4x2 + 16x Find the critical numbers. (Enter your answers as a comma-separated list. If an answer does not exist, enter DNE.) Find the open intervals on which the function is increasing or decreasing. Use a graphing utility to verify your results. (Enter your answers using interval notation. If an answer does not exist, enter DNE.)
Answer #1

- (1) (a) f() 48x -4. Critical points are those points where curre Changes, ito slope. sign of df 48-8X. for critical point d- 1. x = 0, 1, and 1 +49 1.C y = dy 24 an do 20 dx for critical points. 226 no 1.d. ton) = 9 -971. dt -9 < 0 ar so tr, < 0 df